A form of matter is found to have 13 electrons 16 protons and 18 neutrons what is the charge of the matter


1
Expert's answer
2021-07-23T11:01:02-0400

Net charge = protons - electrons

= 16-13 = +3

Charge = 3 × 18 = 54